# SaraCoder: Orchestrating Semantic and Structural Cues for Profit-Oriented Repository-Level Code Completion

### v1

@@ Abstract
Retrieval-augmented generation (RAG) for repository-level code completion commonly relies on superficial text similarity, leading to results plagued by semantic misguidance, redundancy, and homogeneity, while also failing to resolve external symbol ambiguity. To address these challenges, we introduce SaraCoder, a Hierarchical Feature-Optimized retrieval framework. Its core Hierarchical Feature Optimization module systematically refines candidates by distilling deep semantic relationships, pruning exact duplicates, assessing structural similarity with a novel graph-based metric that weighs edits by their topological importance, and reranking results to maximize both relevance and diversity. Furthermore, an External-Aware Identifier Disambiguator module accurately resolves cross-file symbol ambiguity via dependency analysis. Extensive experiments on the challenging CrossCodeEval and RepoEval-Updated benchmarks demonstrate that SaraCoder significantly outperforms existing baselines across multiple programming languages and models. Our work proves that systematically refining retrieval results across multiple dimensions provides a new paradigm for building more accurate and robust repository-level code completion systems.
@@
Introduction
Code Large Language Models (CLLMs) (Izadi, Gismondi, and Gousios 2022; Li et al. 2022b; Allal et al. 2023) built on the Transformer architecture and trained on massive code datasets, have significantly empowered modern software development. They compress vast programming knowledge into billions or trillions of parameters, leading to successful deployments that streamline real-world development and boost productivity. However, as software projects grow in complexity, traditional Code LLMs struggle with repository-level completion. Their limited context windows cannot grasp long-range, cross-file dependencies, preventing a deep understanding of project semantics and architecture. This results in localized, often insufficient, suggestions. Directly processing entire repositories also causes inefficient retrieval, high latency, and an inability to leverage crucial higher-level software engineering knowledge.
To tackle these challenges, Retrieval-Augmented Generation (RAG) (Tang et al. 2023; Zan et al. 2022; Zhang et al. 2023) offers a compelling solution. RAG overcomes the limitations of traditional models by incorporating an external knowledge retrieval mechanism. Within this framework, an efficient retriever locates relevant code snippets, API documentation, or type definitions from the repository in real time. A generation model then integrates these retrieved results with the current context to create syntactically correct, semantically consistent, and standard-compliant code suggestions. RAG is progressively becoming a foundational technology for achieving accurate and trustworthy repository-level intelligent code completion.
Repository-level code completion can be divided into two key scenarios: In-File and Cross-File. In-File completion relies solely on the current file’s context. It is a fundamental and high-frequency task in repository-level intelligence. Cross-File completion involves dependencies on symbols from other files. The true complexity of modern software lies in these inter-module interactions. Cross-File completion is a crucial aspect of repository-level intelligence. However, some existing retrieval-augmented solutions rely too heavily on similarity search without proper filtering, and they often lack the crucial injection of symbolic relationships. This strategy exposes significant flaws in two types of scenarios, as shown in Figure 1: (1) In-File Scenarios: Similarity-based retrieval exhibits two critical flaws in practice. (a) Misleading Superficial Similarity occurs when retrievers recall code fragments that share surface-level characteristics but lack semantic relevance to the current context. The approach suffers from (b) Redundant Retrieval Homogenization, where duplicate or near-identical snippets are repeatedly retrieved, unnecessarily consuming valuable context window capacity that could be allocated for more diverse and informative content. (2) Cross-File Scenarios: Besides the limitations also observed in In-File scenarios, (c) External Symbol Ambiguity is a more severe limitation. Retrievers fail to capture critical dependencies such as unreferenced classes, interface constraints, or architectural rules, triggering error cascades including type inference failures, method signature mismatches, fabricated parameters/attributes, and architecture violations.
@@ Introduction
To address these challenges, we introduce SaraCoder, a Semantic-Aware Code Retrieval framework with Redundancy Reduction and Ambiguity Suppression for Repository-Level Code Completion. SaraCoder implements a hierarchical feature-optimized retrieval-augmented code completion framework, that goes beyond surface-level similarity by extracting deeper relationships such as control and data flow from code snippets. To further refine retrieval quality, we use semantic alignment distillation to capture deep semantic relationships. This is paired with a graph-based structural similarity metric that assesses a candidate’s structural proximity to the target context, weighting edit operations by their topological importance. To combat redundant retriveal homogenization, we integrate deduplication pruning and diversity-aware reranking, ensuring relevance while maximizing diversity. Furthermore, to resolve external symbol ambiguity of Cross-File scenarios in repository-level code completion, we add an external-aware identifier disambiguator that analyzes project-level dependencies. Our key contributions are:
  • We propose SaraCoder, a hierarchical and feature-optimized retrieval-augmented code completion framework. Designed for repository-level code completion, SaraCoder provides a comprehensive solution that prioritizes high-quality, diverse, and streamlined suggestions.
  • SaraCoder seamlessly integrates semantic alignment, redundancy pruning, topological proximity assessment, and diversity-driven reranking, resolving misleading superficial similarity and redundant retrieval homogenization. It also incorporates an external-aware identifier disambiguator that accurately resolves external symbol ambiguities by leveraging project-wide import dependencies.
  • SaraCoder’s coordination of semantic and structural signals enables strong performance even in large, resource-constrained codebases. Its design allows it to complement orthogonally other cross-file methods, delivering synergistic improvements when used in combination.
Related Work
@@
Current RAG methods for repo-level code completion mainly rely on code similarity or cross-file dependencies.
Similar Code Snippet Retrieval for RAG
@@
This approach enhances the quality of code LLM generation by retrieving semantically similar code snippets and integrating them into prompts, mimicking the reference behaviors of programmers. CodeSearchNet (Husain et al. 2020) pioneers large-scale code corpora construction, providing retrieval-based completion references; CodeRetriever (Li et al. 2022a) integrates pretrained models like CodeBERT (Feng et al. 2020) to enhance complex scenario handling; ReACC(Lu et al. 2022) combines vector and string-based retrieval to significantly optimize long-code processing; GraphCoder (Liu et al. 2024) improves code completion by using program dependencies for structured representations, allowing coarse-to-fine retrieval for Python and Java. However, like many other methods, its dependency analysis does not fully grasp deep semantic relationships in code. Additionally, most approaches rely too much on surface-level textual similarity. This often results in redundant retrieved content, wasting resources.
Cross-File Dependency Retrieval Augmentation
@@
This method approaches code completion in complex repositories by leveraging cross-file code context (e.g., dependencies, dataflow, and subsequent similar code). Inspired by Ding et al.’s (Ding et al. 2023) observation that subsequent content of high-similarity snippets effectively informs completion, it injects these snippets into prompts. COCOMIC (Ding et al. 2024) dynamically fuses the context of this file with the cross-file entities retrieved by CCFinder (compressed into [SUM] vectors) through a joint attention mechanism to achieve location-aware code completion. DraCo (Cheng, Wu, and Hu 2024) extends this paradigm through dataflow-guided retrieval, parsing private repositories into code entities and constructing a repository-specific context graph reflecting dependencies. DraCo retrieves precise contextual knowledge from this graph to generate well-structured prompts, overcoming cross-file information barriers and repository-specific accuracy gaps. Current limitations include Python-exclusive implementation with type-sensitive dependencies lacking multilingual support.
Repository-Level Code Completion Evaluation
Although traditional code completion benchmarks (Chen et al. 2021; Austin et al. 2021) focus primarily on local file contexts and isolated code snippets, the inherent complexity of modern software development characterized by vast codebases and intricate in-file dependencies requires a more sophisticated evaluation approach. To address this gap, specialized benchmarks like RepoEval (Zhang et al. 2023), CrossCodeEval (Ding et al. 2023), RepoBench (Liu, Xu, and McAuley 2024), and ReccEval (Cheng, Wu, and Hu 2024) allow for evaluation of repository-level code completion. These establish standardized test scenarios spanning multiple languages, project scales, and contextual complexity levels to ensure rigorous and reproducible assessment. For example, RepoEval targets In-File completion; CrossCodeEval and ReccEval assess Cross-File completion requiring deep contextual understanding.
@@
Method
As shown in Figure 2, SaraCoder is a hierarchical feature-optimized retrieval-enhanced code completion framework, applicable to both In-File and Cross-File completion scenarios. Formally, given a code context $C_{context}=\{x_{1},x_{2},\cdots,x_{n}\}$ and its containing file path $F$, the task aims to predict the next statement $\tilde{y}$.
Database and Initial Candidate Construction
@@
To better represent code logic, we introduce a multi-level code context graph model that integrates control flow, data dependency, and control dependency (Liu et al. 2024). This structured representation offers enhanced generalization capabilities compared to serialization methods, enabling more effective capture of task-relevant context and facilitating easier adaptation to other languages. We utilize program slicing to generate precisely mapped, task-relevant subgraphs from source code on-demand, constructing structured codebases tailored to support specific analysis tasks. When a code completion task request occurs, we extract the Unfinished Code Context $C_{\text{context}}$ and the Import Statements $I$ from the code file . $C_{\text{context}}$ is then used to retrieve an initial candidate set of $top\_k\times p$  [^1] To ensure a high-quality final candidate set of top_ktop\_k results, we expand the initial candidate pool to $top\_k\times p$, allowing more candidates to participate in the Hierarchical Feature Optimization. code snippets $C$ via text similarity from structured codebases.
Hierarchical Feature Optimization (HF_OP)
Semantic Alignment Distillation (SAD)
@@
SAD addresses Superficial Similarity Misguidance by leveraging the GraphcodeBERT (Guo et al. 2021), a pretrained model specialized in code understanding, to capture deep semantic relationships between code snippets. First, the query code $Q$ and candidate set $C$ are tokenized into subword sequences and uniformly padded or truncated to a fixed length $L=512$. Subsequently, during the feature encoding phase, a $768$-dimensional semantic vector $v_{s}$ is extracted for each code unit $s\in{Q}\cup C$, with vector space standardized through L2 normalization. When code repositories lack sufficient repetitive or relevant code, standard filtering methods are too strict, often leading to zero-candidate scenarios. This scarcity of reference material then hurts the accuracy of large language models. To fix this “one-size-fits-all” problem, we introduce a new dynamic quantile threshold mechanism. During the dynamic filtering phase, the cosine similarity set $S=\{\cos(v_{Q},v_{c})\mid c\in C\}$ is computed between the query vector $v_{Q}$ and all candidate vectors $v_{c}$. An adaptive threshold $\tau=\text{quantile}(S,0.75)$ is set at the 75th percentile, outputting filtered results $C_{SAD}=\{c\mid\cos(v_{Q},v_{c})\geq\tau\}$. To reduce redundant computation overhead and improve efficiency, a caching mechanism stores encoding results for high-frequency code.
Redundancy-Aware Pruning (RAP)
@@
This module implements lightweight hash-based deduplication via exact text matching. Using the MD5 algorithm (Rivest 1992), it generates 128-bit hash fingerprints (single computation $\approx 0.02$ms, memory footprint $32$ bytes/hash) to eliminate verbatim duplicates from candidate set $C_{SAD}$ with minimal computational cost, significantly reducing downstream overhead. The module maintains a global hash set $H_{\text{seen}}$ to dynamically track processed sample fingerprints: for each candidate $c\in C_{SAD}$, if its MD5 hash $h_{c}\notin H_{\text{seen}}$, $c$ is added to deduplicated result set $C_{RAP}$ and $H_{\text{seen}}$ is updated. This achieves real-time processing with $O(N)$ time complexity. After SAD processing, the number of code snippets requiring MD5 hashing is limited and their structure is fixed by syntactic and semantic constraints. The MD5 collision resistance (theoretical probability $\approx 1.47\times 10^{-18}$) is sufficient for strict sensitivity. Additionally, MD5’s superior speed and lower memory footprint provide optimal cost-performance.
Topological Proximity Metric (TPM)
@@
At this layer, the decaying subgraph edit distance (D-SED) is introduced to measure the graph similarity between the query graph $G_{q}$ and the candidate graph $G_{c}$ (Ranjan et al.; Zeng et al.).
@@
Diversity-Aware Reranking (DAR)
This module implements a variability-aware ranking model based on the Maximal Marginal Relevance (MMR) (2017) algorithm to maximize result diversity while preserving relevance. It addresses homogeneity in traditional rankings through adversarial similarity calculation and dynamic weight adjustment.
[…]
$S$ contains items $(c,s)\in Q_{TPM}$ that have not been selected into $C_{final}$ yet. $\text{Sim}_{1}$ represents the relevance ($s_{i}=\pi_{2}\circ\iota_{c_{i}}(S)$) of item $c_{i}$ to query $q$. $\text{Sim}_{2}$ denotes the maximum cosine similarity between $c_{i}$ and any item $c_{j}$ in the selected set $C_{final}$. $\lambda$ is a trade-off parameter that balances the emphasis between relevance ($\text{Sim}_{1}$) and diversity ($\text{Sim}_{2}$).
Figure 3: Prompt template used in SaraCoder.
External-Aware Identifier Disambiguator (EAID)

### v2

# SaraCoder: Orchestrating Semantic and Structural Cues for Resource-Optimized Repository-Level Code Completion

@@ Abstract
Despite Retrieval-Augmented Generation improving code completion, traditional retrieval methods struggle with information redundancy and a lack of diversity within limited context windows. To solve this, we propose a resource-optimized retrieval augmentation method, SaraCoder. It maximizes information diversity and representativeness in a limited context window, significantly boosting the accuracy and reliability of repository-level code completion. Its core Hierarchical Feature Optimization module systematically refines candidates by distilling deep semantic relationships, pruning exact duplicates, assessing structural similarity with a novel graph-based metric that weighs edits by their topological importance, and reranking results to maximize both relevance and diversity. Furthermore, an External-Aware Identifier Disambiguator module accurately resolves cross-file symbol ambiguity via dependency analysis. Extensive experiments on the challenging CrossCodeEval and RepoEval-Updated benchmarks demonstrate that SaraCoder outperforms existing baselines across multiple programming languages and models. Our work proves that systematically refining retrieval results across multiple dimensions provides a new paradigm for building more accurate and resource-optimized repository-level code completion systems.
1 Introduction
Code Large Language Models (Code LLMs) Izadi et al. (2022); Li et al. (2022b); Allal et al. (2023), built on the Transformer architecture and trained on massive code corpora. These models compress vast programming knowledge into hundreds of millions of parameters and have been successfully applied in numerous real-world development scenarios, and have significantly advanced the intelligence of modern software development. However, these models often only process local information within their context window. As codebases grow and development tasks become increasingly complex, this limitation becomes more pronounced. Tasks like understanding functionality and fixing bugs often require integrating a wide range of context, such as related API definitions, dependent modules, and type constraints. Since the models cannot perceive information outside their window, the suggestions they generate tend to be localized, often insufficient, and inaccurate.
Retrieval-Augmented Generation (RAG) Tang et al. (2023); Zan et al. (2022); Zhang et al. (2023) overcomes the limitations of traditional models by introducing an external knowledge retrieval mechanism. In this framework, an efficient retriever can dynamically fetch relevant code snippets, API documentation, or type definitions from an external codebase in real time, thereby expanding the model’s contextual awareness. The generative model then integrates these retrieval results with the current context to produce syntactically correct, semantically consistent, and standard-compliant code suggestions. Retrieval-Augmented Generation is gradually becoming a foundational technology for achieving accurate and trustworthy repository-level intelligent code completion.
However, current Retrieval-Augmented Generation (RAG) methods for code completion often have a single-aspect information problem. Methods like GraphCoder Liu et al. (2024b) focus on retrieving similar code snippets, while others like DraCo Cheng et al. (2024) prioritize cross-file context. This narrow focus overlooks the need to combine different types of information, even though developers in real-world settings require both cross-file context for project-wide consistency and similar code snippets for structural reference and efficiency. Limited resources, such as a finite context window, make the effective fusion of these different information types a critical challenge. To make matters worse, existing similarity-based retrieval methods have three major practical flaws. (a) Misleading Superficial Similarity: They often retrieve irrelevant code snippets that share surface-level similarities, which can mislead the model and result in incorrect or ineffective suggestions. (b) Redundant Retrieval Homogenization: Relying solely on similarity ranking frequently yields redundant or duplicate snippets. This wastes valuable context window space and provides a homogeneous, narrow viewpoint, hindering the model’s ability to generate innovative or superior solutions. (c) External Symbol Ambiguity: These methods fail to capture essential dependencies like unreferenced classes or architectural rules. This ambiguity can trigger a cascade of errors, including failed type inference and mismatched signatures, compromising the correctness and reliability of the generated code. Ultimately, these flaws not only degrade the quality of retrieved information but also exacerbate the challenge of a finite context window by filling the limited space with irrelevant, repetitive, or ambiguous data, thereby compromising the generation of correct and innovative code.
@@ 1 Introduction
To address these challenges, we introduce SaraCoder, a Semantic-Aware Code Retrieval framework with Redundancy Reduction and Ambiguity Suppression for Repository-Level Code Completion, that leverages hierarchical resource-optimization to enhance repository-level code completion under resource-constrained conditions. By deeply mining relationships in code snippets, removing redundancy and optimized reranking, we provide LLMs with richer, higher-quality reference cases while minimizing context window consumption. To address misleading superficial similarity, we utilize semantic alignment distillation to capture deep semantic relationships and a graph-based structural similarity metric, which weighs editing operations by topological importance to assess the structural proximity of candidates to the target context. To combat redundant retrieval homogenization, we integrate MD5-based deduplication pruning and diversity-aware reranking, ensuring relevance while maximizing diversity. Additionally, to resolve external symbol ambiguity in repository-level code completion, we introduce an external-aware identifier disambiguator that analyzes project-level dependencies for LLMs. Our key contributions are:
  • We propose SaraCoder, a hierarchical and resource-optimized retrieval-augmented code completion framework. SaraCoder has redefined its code completion goal from retrieving a large volume of relevant code to providing the most valuable information within a limited context.
  • As an intelligent information filter, SaraCoder upgrades code completion by moving from surface-level code matching to intelligent decision-making based on deep semantics and project structure. This provides LLMs with high-quality, precise information, leading to more efficient repository-level code completion within limited contexts.
  • SaraCoder’s design allows it to maintain suggestion quality even with a nearly full context window. This ensures that when complementing other methods that provide different information, it minimizes negative interference and maximizes the preservation of their content. This synergy enables it to complement orthogonally other methods that provide cross-file context, providing a cooperative improvements when used in combination.
2 Related Work
@@
Current Retrieval-Augmented Generation methods for repo-level code completion mainly rely on code similarity or cross-file dependencies.
2.1 Similar Code Snippet Retrieval for RAG
@@
This approach enhances the quality of code LLM generation by retrieving semantically similar code snippets and integrating them into prompts, mimicking the reference behaviors of programmers. CodeSearchNet Husain et al. (2020) pioneers large-scale code corpora construction, providing retrieval-based completion references; CodeRetriever Li et al. (2022a) integrates pretrained models like CodeBERT Feng et al. (2020) to enhance complex scenario handling; ReACC Lu et al. (2022) combines vector and string-based retrieval to significantly optimize long-code processing; GraphCoder Liu et al. (2024b) improves code completion by using program dependencies for structured representations, allowing coarse-to-fine retrieval for Python and Java. However, like many other methods, its dependency analysis does not fully grasp deep semantic relationships in code. Additionally, most approaches rely too much on surface-level textual similarity. This often results in redundant retrieved content, wasting resources.
2.2 Cross-File Dependency Retrieval Augmentation
@@
This method approaches code completion in complex repositories by leveraging cross-file code context (e.g., dependencies, dataflow, and subsequent similar code). Inspired by Ding et al.’s Ding et al. (2023) observation that subsequent content of high-similarity snippets effectively informs completion, it injects these snippets into prompts. COCOMIC Ding et al. (2024) dynamically fuses the context of this file with the cross-file entities retrieved by CCFinder (compressed into [SUM] vectors) through a joint attention mechanism to achieve location-aware code completion. DraCo Cheng et al. (2024) extends this paradigm through dataflow-guided retrieval, parsing private repositories into code entities and constructing a repository-specific context graph reflecting dependencies. DraCo retrieves precise contextual knowledge from this graph to generate well-structured prompts, overcoming cross-file information barriers and repository-specific accuracy gaps. Current limitations include Python-exclusive implementation with type-sensitive dependencies lacking multilingual support.
2.3 Repository-Level Code Completion Evaluation
Traditional code completion benchmarks like Chen et al. (2021); Austin et al. (2021) focus on isolated snippets, but modern software development’s complexity demands a better evaluation. To address this, specialized benchmarks such as RepoEval Zhang et al. (2023), CrossCodeEval Ding et al. (2023), RepoBench Liu et al. (2024a), and ReccEval Cheng et al. (2024) have emerged. They provide standardized, rigorous tests across various languages and project scales. These benchmarks divide repository-level code completion into two scenarios. (1) In-File Completion: A high-frequency task that uses only the current file’s context (e.g., RepoEval). (2) Cross-File Completion: A more complex task that requires understanding and completing code with dependencies on symbols from other files (e.g., CrossCodeEval and ReccEval). This shift highlights the move from simple, isolated tests to comprehensive evaluations that reflect real-world coding environments.
3 Method
As shown in Figure 2, SaraCoder is a hierarchical feature-optimized retrieval-enhanced code completion framework. Formally, given a code context $C_{context}=\{x_{1},x_{2},\cdots,x_{n}\}$ and its containing file path $F$, the task aims to predict the next statement $\tilde{y}$.
@@ 3 Method
3.1 Database and Initial Candidate Construction
@@
To better represent code logic, we introduce a multi-level code context graph model that integrates control flow, data dependency, and control dependency Liu et al. (2024b). This structured representation offers enhanced generalization capabilities compared to serialization methods, enabling more effective capture of task-relevant context and facilitating easier adaptation to other languages. We utilize program slicing to generate precisely mapped, task-relevant subgraphs from source code on-demand, constructing structured codebases tailored to support specific analysis tasks. When a code completion task request occurs, we extract the Unfinished Code Context $C_{\text{context}}$ and the Import Statements $I$ from the code file . $C_{\text{context}}$ is then used to retrieve an initial candidate set of $top\_k\times p$  [^1] To ensure a high-quality final candidate set of top_ktop\_k results, we expand the initial candidate pool to $top\_k\times p$, allowing more candidates to participate in the Hierarchical Feature Optimization. code snippets $C$ via text similarity from structured codebases.
3.2 Hierarchical Feature Optimization (HF_OP)
3.2.1 Semantic Alignment Distillation
@@
Semantic alignment distillation addresses Superficial Similarity Misguidance by leveraging the GraphcodeBERT Guo et al. (2021), a pretrained model specialized in code understanding, to capture deep semantic relationships between code snippets. First, the query code $Q$ and candidate set $C$ are tokenized into subword sequences and uniformly padded or truncated to a fixed length $L=512$. Subsequently, during the feature encoding phase, a $768$-dimensional semantic vector $v_{s}$ is extracted for each code unit $s\in{Q}\cup C$, with vector space standardized through L2 normalization. When code repositories lack sufficient repetitive or relevant code, standard filtering methods are too strict, often leading to zero-candidate scenarios. This scarcity of reference material then hurts the accuracy of large language models. To fix this “one-size-fits-all” problem, we introduce a new dynamic quantile threshold mechanism. During the dynamic filtering phase, the cosine similarity set $S=\{\cos(v_{Q},v_{c})\mid c\in C\}$ is computed between the query vector $v_{Q}$ and all candidate vectors $v_{c}$. An adaptive threshold $\tau=\text{quantile}(S,0.75)$ is set at the 75th percentile, outputting filtered results $C_{SAD}=\{c\mid\cos(v_{Q},v_{c})\geq\tau\}$. To reduce redundant computation overhead and improve efficiency, a caching mechanism stores encoding results for high-frequency code.
3.2.2 Redundancy-Aware Pruning
@@
This module implements lightweight hash-based deduplication via exact text matching. Using the MD5 algorithm Rivest (1992), it generates 128-bit hash fingerprints (single computation $\approx 0.02$ms, memory footprint $32$ bytes/hash) to eliminate verbatim duplicates from candidate set $C_{SAD}$ with minimal computational cost, significantly reducing downstream overhead. The module maintains a global hash set $H_{\text{seen}}$ to dynamically track processed sample fingerprints: for each candidate $c\in C_{SAD}$, if its MD5 hash $h_{c}\notin H_{\text{seen}}$, $c$ is added to deduplicated result set $C_{RAP}$ and $H_{\text{seen}}$ is updated. This achieves real-time processing with $O(N)$ time complexity. After semantic alignment distillation processing, the number of code snippets requiring MD5 hashing is limited and their structure is fixed by syntactic and semantic constraints. The MD5 collision resistance (theoretical probability $\approx 1.47\times 10^{-18}$) is sufficient for strict sensitivity. Additionally, MD5’s superior speed and lower memory footprint provide optimal cost-performance.
3.2.3 Topological Proximity Metric
@@
At this layer, the decaying subgraph edit distance (D-SED) is introduced to measure the graph similarity between the query graph $G_{q}$ and the candidate graph $G_{c}$ (Ranjan et al.; Zeng et al.). A higher D-SED value indicates less similarity. We calculate D-SED for code snippets to quantify their structural similarity and retain those with the closest match.
@@
3.2.4 Diversity-Aware Reranking
This module implements a variability-aware ranking model based on the Maximal Marginal Relevance (MMR) Carbonell & Goldstein (2017) algorithm to maximize result diversity while preserving relevance. It addresses homogeneity in traditional rankings through adversarial similarity calculation and dynamic weight adjustment. $S$ contains items $(c,s)\in Q_{TPM}$ that have not been selected into $C_{final}$ yet. $\text{Sim}_{1}$ represents the relevance ($s_{i}=\pi_{2}\circ\iota_{c_{i}}(S)$) of item $c_{i}$ to query $q$. $\text{Sim}_{2}$ denotes the maximum cosine similarity between $c_{i}$ and any item $c_{j}$ in the selected set $C_{final}$. $\lambda$ is a trade-off parameter that balances the emphasis between relevance ($\text{Sim}_{1}$) and diversity ($\text{Sim}_{2}$).
$\text{MMR}=\arg\max_{c_{i}\in S}\Bigl[\lambda\cdot\text{Sim}_{1}(c_{i},q)-(1-\lambda)\cdot\max_{c_{j}\in C_{\text{final}}}\text{Sim}_{2}(c_{i},c_{j})\Bigr]$ (2)
3.3 External-Aware Identifier Disambiguator (EAID)